# Fractional Semilinear Equations on Hyperbolic Spaces

### v1

@@ 4.1. Subcritical case $1<p<\frac{n+2s}{n-2s}$
Fix $\varepsilon>0$. Choose $R>0$ large and write
$u(x)=\int_{B_{R}(o)}G(\rho(x,y))u(y)^{p}\,dV_{y}+\int_{\mathbb{H}^{n}\setminus B_{R}(o)}G(\rho(x,y))u(y)^{p}\,dV_{y}.$
For the local part, we have that for any fixed $y\in B_{R}$, $\rho(x,y)\to\infty$ as $\rho(x,o)\to\infty$. By kernel estimate, $G(\rho(x,y))\lesssim\rho^{-1+s}e^{-(n-2s)\rho(x,o)}$. Hence
$\int_{B_{R}}G(\rho(x,y))u(y)^{p}\,dV_{y}\lesssim\rho^{-1+s}e^{-(n-2s)\rho(x,o)}\int_{B_{R}}u(y)^{p}\,dV_{y}\to 0$
uniformly in $x$ as $\rho(x,o)\to\infty$. ∎
We then establish the following result regarding the improvement of integrability.
Proposition 4.3.
If $u\in L^{p+1}$ is a solution of (1.7) with $1<p<\frac{n+2s}{n-2s}$, then $u\in L^{\infty}(\mathbb{H}^{n})$.
Proof.
We use the equivalent integral equation of (1.7),
$u(x)=\int_{H^{n}}G(\rho(x,y))\,u(y)^{p}\,dV_{y}.$
Since $u\in L^{r}$ we have $u^{p}\in L^{a_{1}}$ with $a:=\frac{r}{p}<\frac{2n}{n+2s}<\frac{n}{2s}$, the HLS inequality implies that
$u=I_{2s}(u^{p})\in L^{t_{1}}(\mathbb{H}^{n}),\qquad\frac{1}{t_{1}}=\frac{p}{r}-\frac{2s}{n}.$
Direct calculation verifies $t_{1}>r$ for any $p<\frac{n+2s}{n-2s}$. Again, $u^{p}\in L^{a_{2}}$ with $a_{2}:=\frac{t_{1}}{p}<1<\frac{n}{2s}$, and HLS inequality implies
$u=I_{2s}(u^{p})\in L^{t_{2}}(\mathbb{H}^{n}),\qquad\frac{1}{t_{2}}=\frac{p}{t_{1}}-\frac{2s}{n}.$
Thus,
$\frac{1}{t_{k+1}}=\frac{p}{t_{k}}-\frac{2s}{n},\quad k=0,1,2,\ldots,\text{ and }t_{0}=r$
or equivalently,
$t_{k+1}=\frac{1}{p/t_{k}-2s/n\,}.$
We choose $t_{0}=p+1<\frac{2n}{n-2s}$ so that $t_{1}=\frac{n(p+1)}{np-2s(p+1)}>p+1$. Direct calculation shows that $t_{k}$ is increasing in $k$ when $t_{k}\in(a,b)$, where $a=\frac{n(p-1)}{2s}$ and $b=\frac{np}{2s}$. Moreover, if we denote $x_{k}=\frac{1}{t_{k}}$, we obtain
$x_{k}=p^{k}\left(\frac{1}{t_{0}}-\frac{1}{a}\right)+\frac{1}{a}.$
Define $K:=\min\{k:\ x_{k}<\frac{1}{b}\},$ we solve the above equation and get $K\;=\;\left\lceil\log_{p}\Big(\frac{1/t_{0}-1/b}{1/t_{0}-1/t_{0}}\Big)\right\rceil$, where $\lceil\,\rceil$ is the ceiling function. Thus, we can apply the HLS inequality for $k=0,1,2,\ldots,K-1$ to obtain $u\in L^{t_{K}}$ with $t_{K}>\frac{np}{2s}$. Then, applying Holder’s inequality to the integral equation, we have
$u(x)$ $=\int G(x,y)\,u(y)^{p}\,dy\lesssim{\left\lVert{G}\right\rVert}_{L^{a^{\prime}}}{\left\lVert{u^{p}}\right\rVert}_{L^{a}}$
$={\left\lVert{G}\right\rVert}_{L^{a^{\prime}}}{\left\lVert{u}\right\rVert}_{L^{ap}}^{p},$
where $\frac{1}{a}+\frac{1}{a^{\prime}}=1,ap>\frac{np}{2s}.$ Since
$\|G(x,\cdot)\|_{L^{a^{\prime}}}^{a^{\prime}}=\int_{\mathbb{H}^{n}}G(\rho(x,y))^{a^{\prime}}\,dV_{y}=\omega_{n-1}\int_{0}^{\infty}K_{s}(\rho)^{a^{\prime}}\sinh^{n-1}\rho\,d\rho.$
Using the asymptotic behavior of $K_{s}(\rho)\sim C\rho^{-(n-2s)}$ when $\rho\to 0^{+}$, we see that
$\int_{0}^{\infty}K_{s}(\rho)^{a^{\prime}}\sinh^{n-1}\rho\,d\rho\sim$ $\int_{0}^{r}\rho^{-a^{\prime}(n+2s)}\rho^{n-1}\,d\rho+\int_{1}^{\infty}e^{-a^{\prime}(n-1)\rho}\,e^{(n-1)\rho}\,d\rho$
$=$ $\int_{0}^{1}\rho^{n-1-a^{\prime}(n+2s)}\,d\rho+\int_{1}^{\infty}e^{-(a^{\prime}(n-1)-(n-1))\rho}\,d\rho$
is finite provided that $n-a^{\prime}(n+2s)<0$, or equivalently $ap<\frac{np}{2s}$. Thus,
${\left\lVert{u}\right\rVert}_{L^{\infty}(\mathbb{H}^{n})}=\sup_{\mathbb{H}^{n}}\int_{\mathbb{H}^{n}}G(x,y)\,u(y)^{p}\,dy\lesssim\|u\|_{L^{r}}^{p}<\infty,$
and the lemma is proved.
@@
For $1<p<\frac{n+2k}{n-2k}$, we show that (1.7) admits no positive solution via a moving plane method.
Recall the foilation structure of $\mathbb{H}^{n}$ and the notations:
@@
$x_{\lambda}=I_{\lambda}(x),\quad\lambda\in\mathbb{R}.$
@@
$w_{\lambda}(x)=u(x_{\lambda})-u(x).$ (4.4)
Then $w_{\lambda}$ satisfies the following equation
$(-\Delta_{\mathbb{H}})^{s}w_{\lambda}(x)=u(x_{\lambda})^{p}-u(x)^{p}>pw_{\lambda}u(x)^{p}\text{ in }\Sigma^{-}_{\lambda}.$
Denote
$c_{\lambda}(x):=-pu(x)^{p},$
then we have
$(-\Delta)^{s}w_{\lambda}+c_{\lambda}(x)w_{\lambda}(x)\geqslant 0.$
@@
$\Sigma_{t}^{-}:=\{x\in\Omega_{t}:w_{t}(x):=u(x^{t})-u(x)<0\}.$
We want to prove for $t$ sufficiently negative, $\Sigma_{t}^{-}=\emptyset$. We first have
@@
$=\int_{\Sigma_{\lambda}}G(x,y)u^{p}(y)dV_{y}+\int_{\Sigma^{c}_{\lambda}}G(x,y)u^{p}(y)dV_{y}$
$\quad-\int_{\Sigma_{\lambda}}G(x,y)u^{p}(y_{\lambda})dV_{y}-\int_{\Sigma^{c}_{\lambda}}G(x,y)u^{p}(y_{\lambda})dV_{y}$
$=\int_{\Sigma_{\lambda}}G(x,y)u^{p}(y)dV_{y}+\int_{\Sigma_{\lambda}}G(x,y_{\lambda})u^{p}(y_{\lambda})dV_{y}$
$\quad-\int_{\Sigma_{\lambda}}G(x,y)u^{p}(y_{\lambda})dV_{y}-\int_{\Sigma_{\lambda}}G(x,y_{\lambda})u^{p}(y)dV_{y}$
$=\int_{\Sigma_{\lambda}}G(x,y)u^{p}(y)dV_{y}+\int_{\Sigma_{\lambda}}G(x_{\lambda},y)u^{p}(y_{\lambda})dV_{y}$
$\quad-\int_{\Sigma_{\lambda}}G(x,y)u^{p}(y_{\lambda})dV_{y}-\int_{\Sigma_{\lambda}}G(x_{\lambda},y)u^{p}(y)dV_{y}$
$=\int_{\Sigma_{\lambda}}\left(G(x,y)-G(\bar{x},y)\right)\left(u^{q}(y)-u^{q}(y_{\lambda})\right)dV_{y}$
Hence, by the definition of $\Sigma_{\lambda}^{-}$ and the monotonicity of $G$, we have
$u(x)-u_{\lambda}(x)$ $\leqslant\int_{\Sigma_{\lambda}^{-}}G(x,y)\left(u^{q}(y)-u^{q}(\bar{y}))\right)dV_{y}.$
Fix a radius $R>0$, for any $x,y\in\mathbb{H}^{n}$, we decompose
$G(x,y)=G_{1}(x,y)+G_{2}(x,y),$
@@
$I_{i}(x):=\int_{\mathbb{H}^{n}}G_{i}(x,y)\,u(y)^{p}\,d\mu(y),\qquad i=\{1,2\}.$
For near part, we observe that in any fixed geodesic ball of radius $R$, the hyperbolic metric is equivalent to the Euclidean metric. Hence the classical HLS inequality applies locally, and one obtains the estimate
$\|I_{1}\|_{L^{q}(\Sigma_{t}^{-})}\leq C_{R}\,\|u^{p}\chi_{\Sigma_{t}^{-}}\|_{L^{q^{\prime}}(\Omega_{t})}$ (4.5)
which holds provided $1/q=1/q^{\prime}-2s/n$. The constant $C_{R}$ depends on $R$ and on the chart distortion but is finite for fixed $R$.
For the far part, use the exponential decay of $G$. We have $G(x,y)\leq Cd^{-1+s}e^{-(n-1)d(x,y)}$ for large distances. Then for $x\in\Omega_{t}$,
$|I_{2}(x)|\leq\int_{d(x,y)>R}G(x,y)u(y)^{p}\,d\mu(y).$
Using Hölder’s inequality, we have for $1/s+1/q^{\prime}=1$,
$|I_{2}(x)|\leqslant{\left\lVert{G(x,\cdot)}\right\rVert}_{L^{s}(\{d(x,y)>R\})}{\left\lVert{u^{p}}\right\rVert}_{L^{q^{\prime}}(\{d(x,y)>R\})}$
@@
$\|I_{2}\|_{L^{q}(\Sigma_{t}^{-})}$ $\leq\Big(\int_{\Sigma_{t}^{-}}\|G(x,\cdot)\|_{L^{s}(\mathbb{H}^{n}\setminus B_{R})}^{q}\,d\mu(x)\Big)^{1/q}\;\cdot\;\|u^{p}\chi_{\Omega_{t}}\|_{L^{q^{\prime}}(\Omega_{t})}$
$\leq\sup_{x\in\Sigma_{t}^{-}}\|G(x,\cdot)\|_{L^{s}(\mathbb{H}^{n}\setminus B_{R})}\;\big(\mu(\Sigma_{t}^{-})\big)^{1/q}\ \|u^{p}\chi_{\Omega_{t}}\|_{L^{q^{\prime}}}.$ (4.6)
@@
$\|w_{t}\|_{L^{q}(\Sigma_{t}^{-})}$ $\leqslant C_{R}\|u^{p}\|_{L^{q^{\prime}}(\chi_{\Sigma_{t}^{-}})}+\varepsilon(R,t)\,\big(\mu(\Sigma_{t}^{-})\big)^{1/q}\|u^{p}\|_{L^{q^{\prime}}(\chi_{\Sigma_{t}^{-}})}$
$\lesssim\big(\mu(\Sigma_{t}^{-})\big)^{1/q}\|u^{p}\|_{L^{q^{\prime}}(\chi_{\Sigma_{t}^{-}})}.$
@@
$\|w_{t}\|_{L^{q}(\Sigma_{t}^{-})}$ $\leqslant\big(\mu(\Sigma_{t}^{-})\big)^{1/q}{\left\lVert{u_{\lambda}^{p}-u^{p}}\right\rVert}_{L^{q^{\prime}}(\Sigma_{t}^{-})}\leqslant\big(\mu(\Sigma_{t}^{-})\big)^{1/q}{\left\lVert{p\xi^{p-1}(u_{\lambda}-u)}\right\rVert}_{L^{q^{\prime}}(\Sigma_{t}^{-})}$
$\leqslant\big(\mu(\Sigma_{t}^{-})\big)^{1/q}{\left\lVert{u^{p-1}}\right\rVert}_{L^{n/2s}(\Sigma_{t}^{-})}{\left\lVert{u_{\lambda}-u}\right\rVert}_{L^{q}(\Sigma_{t}^{-})}.$
@@
When $\varepsilon$ is small, $\Sigma_{\lambda}^{-}$ is close to zero so that $\|u_{t}-u\|_{L^{q}(\Sigma_{\lambda}^{-})}=0$. This imples that we can keep moving the plane $U_{\lambda_{0}}$. Now we see $u(x)\leqslant u(\bar{x})$ with respect to $\Sigma_{\lambda_{0}}$. A similar argument shows $u(\bar{x})\geqslant u(x)$ by rotation. Consequently, there exsits a point $p$ such that $u$ is constant on the geodesic spheres center at $p$.
Since $u\in L^{\infty}\cap L^{p+1}$, we obtain the following due to the bottom of the spectrum of $\Delta_{\mathbb{H}}$ on $\mathbb{H}^{n}$ being $\lambda_{0}=\frac{(n-1)^{2}}{4}$:
$\langle u,(-\Delta)^{s}u\rangle_{L^{2}(\mathbb{H}^{n})}=\int_{\mathbb{H}^{n}}u^{p+1}\,dV\geqslant\lambda_{0}^{s}\|u\|_{L^{2}}^{2}.$
Sobolev inequality in hyperbolic space gives
$\|u\|_{L^{p+1}}\leqslant C\|(-\Delta)^{s/2}u\|_{L^{2}}=C\langle u,(-\Delta)^{s}u\rangle^{1/2}.$
On the other hand, we have
$\langle u,(-\Delta)^{s}u\rangle_{L^{2}(\mathbb{H}^{n})}=\int_{\mathbb{H}^{n}}u^{p+1}\,dV.$
Therefore
$\|u\|_{L^{p+1}}^{p+1}\leq C^{p+1}\left(\|u\|_{L^{p+1}}^{p+1}\right)^{(p+1)/2}$
Let $X:=\|u\|_{L^{p+1}}^{p+1}$, then
$X\leq C^{p+1}X^{(p+1)/2}.$
Divide both sides by $X^{\frac{p+1}{2}}$, we obtain $X^{\frac{1-p}{2}}\leq C^{p+1}$. However, this is true only if $X=0$, due to scaling invariance. To see this, wefine the rescaled family $u_{\lambda}(|x|)=\lambda^{\alpha}u(\lambda|x|)$, and $X_{\lambda}:=\|u_{\lambda}\|_{L^{p+1}}^{p+1}=\lambda^{\beta}X$. Since $\beta\neq 0$, by choosing $\lambda\to 0$ or $\lambda\to\infty$, we can make $X_{\lambda}\to 0^{+}$, so that $X_{\lambda}^{\frac{1-p}{2}}\to+\infty$ since $p>1$.

### v2

@@ 4.1. Subcritical case $1<p<\frac{n+2s}{n-2s}$
If $u\in H^{s}$ is a weak solution of (1.8) with $1<p<\frac{n+2s}{n-2s}$, then $u\in L^{\infty}(\mathbb{H}^{n})$.
We then establish the decay property of $u$ at infinity.
Lemma 4.3.
@@
Fix $\varepsilon>0$. Choose $\delta>0$, and $R>0$ large and split the integral into three parts:
$\mathbb{H}^{n}=B_{R}(o)\cup(\mathbb{H}^{n}\setminus B_{R}(o))\cap B_{\delta}(x)\cup(\mathbb{H}^{n}\setminus B_{R}(o))\setminus B_{\delta}(x).$
For the local part, we have that for any fixed $y\in B_{R}$, $\rho(x,y)\to\infty$ as $\rho(x,o)\to\infty$. By kernel estimate, $G_{s}(\rho(x,y))\lesssim\rho^{-1+s}e^{-(n-1)\rho(x,y)}$. Hence
$\int_{B_{R}(o)}G_{s}(\rho(x,y))u(y)^{p}\,dV_{y}\lesssim\rho(x,o)^{-1+s}e^{-(n-1)(\rho(x,o)-R)}\int_{B_{R}}u(y)^{p}\,dV_{y}<\varepsilon$
as $\rho(x,o)\to\infty$.
For the second part, we have
$\int_{(\mathbb{H}^{n}\setminus B_{R}(o))\cap B_{\delta}(x)}G_{s}(\rho(x,y))u(y)^{p}\,dV_{y}\leqslant{\left\lVert{u}\right\rVert}_{L^{\infty}}^{p}\int_{B_{\delta}(x)}G_{s}(\rho(x,y))\,dV_{y}$
$\lesssim{\left\lVert{u}\right\rVert}_{L^{\infty}}^{p}\int_{0}^{\delta}\rho^{-n+2s}\sinh^{n-1}\rho\,d\rho\lesssim{\left\lVert{u}\right\rVert}_{L^{\infty}}^{p}\int_{0}^{\delta}\rho^{2s-1}\,d\rho<\varepsilon.$
For the last part, using Hölder’s inequality, we have
$\int_{(\mathbb{H}^{n}\setminus B_{R}(o))\setminus B_{\delta}(x)}G_{s}(\rho(x,y))u(y)^{p}\,dV_{y}\lesssim{\left\lVert{G_{s}}\right\rVert}_{L^{a^{\prime}}}{\left\lVert{u^{p}}\right\rVert}_{L^{a}}.$
where $a=\frac{2n}{n+2s}$ and $a^{\prime}=\frac{2n}{n-2s}$. We see that
$\|G_{s}(\rho(x,\cdot))\|_{L^{a^{\prime}}}^{a^{\prime}}\lesssim\omega_{n-1}\int_{\delta}^{\infty}\rho^{-a^{\prime}(-1+s)}e^{-a^{\prime}(n-1)\rho}\sinh^{n-1}\rho\,d\rho<\infty,$
and $\|u^{p}\|_{L^{a}(\mathbb{H}^{n}\setminus B_{R}(o))}<\varepsilon$ for $R$ large enough. Therefore, $u(x)=\int_{\mathbb{H}^{n}}G_{s}(x,y)u(y)^{p}\,dV_{y}\to 0$ as $\rho(x,o)\to\infty$.
@@
Recall the foliation structure of $\mathbb{H}^{n}$ and the notations:
@@
$x_{\lambda}=I_{\lambda}(x),\quad u_{\lambda}(x)=u(x_{\lambda}),\quad\lambda\in\mathbb{R}.$
@@
$w_{\lambda}(x)=u_{\lambda}(x)-u(x).$ (4.6)
@@
$\Sigma_{\lambda}^{-}:=\{x\in\Sigma_{\lambda}:w_{\lambda}(x):=u_{\lambda}(x)-u(x)<0\}.$
We want to prove for $\lambda$ sufficiently negative, $\Sigma_{\lambda}^{-}=\emptyset$. We first have
@@
$=\int_{\Sigma_{\lambda}}G_{s}(x,y)u^{p}(y)dV_{y}+\int_{\Sigma^{c}_{\lambda}}G_{s}(x,y)u^{p}(y)dV_{y}$
$\quad-\int_{\Sigma_{\lambda}}G_{s}(x,y)u^{p}(y_{\lambda})dV_{y}-\int_{\Sigma^{c}_{\lambda}}G_{s}(x,y)u^{p}(y_{\lambda})dV_{y}$
$=\int_{\Sigma_{\lambda}}G_{s}(x,y)u^{p}(y)dV_{y}+\int_{\Sigma_{\lambda}}G_{s}(x,y_{\lambda})u^{p}(y_{\lambda})dV_{y}$
$\quad-\int_{\Sigma_{\lambda}}G_{s}(x,y)u^{p}(y_{\lambda})dV_{y}-\int_{\Sigma_{\lambda}}G_{s}(x,y_{\lambda})u^{p}(y)dV_{y}$
$=\int_{\Sigma_{\lambda}}G_{s}(x,y)u^{p}(y)dV_{y}+\int_{\Sigma_{\lambda}}G_{s}(x_{\lambda},y)u^{p}(y_{\lambda})dV_{y}$
$\quad-\int_{\Sigma_{\lambda}}G_{s}(x,y)u^{p}(y_{\lambda})dV_{y}-\int_{\Sigma_{\lambda}}G_{s}(x_{\lambda},y)u^{p}(y)dV_{y}$
$=\int_{\Sigma_{\lambda}}\left(G_{s}(x,y)-G_{s}(x_{\lambda},y)\right)\left(u^{p}(y)-u^{p}(y_{\lambda})\right)dV_{y}.$
Hence, by the definition of $\Sigma_{\lambda}^{-}$ and the monotonicity of $G_{s}$, we have
$u(x)-u_{\lambda}(x)$ $\leqslant\int_{\Sigma_{\lambda}^{-}}G_{s}(x,y)\left(u^{p}(y)-u^{p}(y_{\lambda})\right)dV_{y},$
since $\rho(x,y)>\rho(x_{\lambda},y)$ for $y\in\Sigma_{\lambda}$. For a fixed radius $R>0$, and any $x,y\in\mathbb{H}^{n}$, we decompose
$G_{s}(x,y)=G_{1}(x,y)+G_{2}(x,y),$
@@
$I_{i}(x):=\int_{\mathbb{H}^{n}}G_{s}^{i}(x,y)\,u(y)^{p}\,d\mu(y),\qquad i=\{1,2\}.$
For near part, we observe that in any fixed geodesic ball of radius $R$, the hyperbolic metric is equivalent to the Euclidean one. Hence the classical HLS inequality applies locally, and one obtains
$\|I_{1}\|_{L^{q}(\Sigma_{\lambda}^{-})}\leq C_{R}\,\|u^{p}\chi_{\Sigma_{\lambda}^{-}}\|_{L^{q^{\prime}}(\Sigma_{\lambda})},$ (4.7)
which holds for $1/q=1/q^{\prime}-2s/n$. For the far part, use the exponential decay of $G_{s}$. We have $G_{s}(x,y)\lesssim d^{-1+s}e^{-(n-1)\rho(x,y)}$. Then for $x\in\Sigma_{\lambda}$,
$|I_{2}(x)|\leq\int_{\rho(x,y)>R}G_{s}(x,y)u(y)^{p}\,d\mu(y).$
Using Hölder’s inequality, we have for $1/r+1/q^{\prime}=1$,
$|I_{2}(x)|\leqslant{\left\lVert{G_{s}(x,\cdot)}\right\rVert}_{L^{r}(\{\rho(x,y)>R\})}{\left\lVert{u^{p}}\right\rVert}_{L^{q^{\prime}}(\{\rho(x,y)>R\})}$
@@
$\|I_{2}\|_{L^{q}(\Sigma_{\lambda}^{-})}$ $\leq\Big(\int_{\Sigma_{\lambda}^{-}}\|G_{s}(x,\cdot)\|_{L^{r}(\mathbb{H}^{n}\setminus B_{R})}^{q}\,dV_{x}\Big)^{1/q}\;\cdot\;\|u^{p}\chi_{\Sigma_{\lambda}}\|_{L^{q^{\prime}}(\Sigma_{\lambda})}$
$\leq\sup_{x\in\Sigma_{\lambda}^{-}}\|G_{s}(x,\cdot)\|_{L^{r}(\mathbb{H}^{n}\setminus B_{R})}\;\big(\mu(\Sigma_{\lambda}^{-})\big)^{1/q}\ \|u^{p}\chi_{\Sigma_{\lambda}}\|_{L^{q^{\prime}}(\Sigma_{\lambda})}.$ (4.8)
@@
$\|w_{\lambda}\|_{L^{q}(\Sigma_{\lambda}^{-})}$ $\leqslant C_{R}\|u^{p}\|_{L^{q^{\prime}}(\chi_{\Sigma_{\lambda}^{-}})}+\varepsilon(R,\lambda)\,\big(\mu(\Sigma_{\lambda}^{-})\big)^{1/q}\|u^{p}\|_{L^{q^{\prime}}(\chi_{\Sigma_{\lambda}^{-}})}$
$\lesssim\big(\mu(\Sigma_{\lambda}^{-})\big)^{1/q}\|u^{p}\|_{L^{q^{\prime}}(\chi_{\Sigma_{\lambda}^{-}})}.$
@@
$\|w_{\lambda}\|_{L^{q}(\Sigma_{\lambda}^{-})}$ $\leqslant\big(\mu(\Sigma_{\lambda}^{-})\big)^{1/q}{\left\lVert{u_{\lambda}^{p}-u^{p}}\right\rVert}_{L^{q^{\prime}}(\Sigma_{\lambda}^{-})}\leqslant\big(\mu(\Sigma_{\lambda}^{-})\big)^{1/q}{\left\lVert{p\xi^{p-1}(u_{\lambda}-u)}\right\rVert}_{L^{q^{\prime}}(\Sigma_{\lambda}^{-})}$
$\leqslant\big(\mu(\Sigma_{\lambda}^{-})\big)^{1/q}{\left\lVert{u^{p-1}}\right\rVert}_{L^{n/2r}(\Sigma_{\lambda}^{-})}{\left\lVert{u_{\lambda}-u}\right\rVert}_{L^{q}(\Sigma_{\lambda}^{-})}.$
@@
When $\varepsilon$ is small, $\Sigma_{\lambda}^{-}$ is close to zero so that $\|u_{t}-u\|_{L^{q}(\Sigma_{\lambda}^{-})}=0$. This implies that we can keep moving the plane $U_{\lambda_{0}}$. Now we see $u(x)\leqslant u(\bar{x})$ with respect to $\Sigma_{\lambda_{0}}$. A similar argument shows $u(\bar{x})\geqslant u(x)$ by rotation. Consequently, there exists a point $P$ such that $u$ is constant on the geodesic spheres center at $P$.